An electron travels 0.01 m across an electric potential difference of 6 x 10^6 V. What is the force exerted on the electron?
1
Expert's answer
2020-04-01T10:15:11-0400

Electric field ="\frac{potential \ difference (\Delta V)}{\Delta x}=\frac{6\times10^6}{0.01}=6\times10^8V/m"


Force ="q\times electric \ field =1.6\times10^{-19}\times6\times10^8=9.6\times10^{-11}N"